Section web-meeting

→ Europe/Zurich

***
# Minutes of the meeting: Teleworking
###### Date: Wednesday, 27 May 2020
###### event Url: https://indico.cern.ch/category/11506/
###### event id: 900940
***

## Participants

Marta, Bernardo, Stathis, Ioannis, Michal, Ruben, 
Salvador, Franco, Yacine, Patrick, Paula, 
Jerome, Gaelle, Gerard, Maryline

**HEALTH STATUS: Everyone is fine without any COVID-19 symptom!**


## Main points 

### Still valid:
* **TW general rules:**
**A**. 1/day: check email
**B**. Be on standby (!), while being easily reachable by phone or email.

* **Quarantine**: We are reaching the T0 and STAGE 1 where CERN will open gradually its gates on 18/5, while the SM18 on 11/8. However, this will change dynamically based on the circumstances.

* **Priorities and staff:**
    - For the moment (started on 27/4) there are 3 operator positions with 100% work allocation at SM18. J. Hurte (Demo2), V. Desbiolles (Potaims - QXFP1), O. Ditsch (Quench antennas - QXFP1). More have been added (B. Mouches (TE12) and A. Azarkan (TE14) on QXFP1 and C1 bench connection at SM18). 
    - In total four operators and two staff will be needed for the tests in this first stage.
    - MBHA001 has the highest priority, with the Demo2, MQXFP1, and MBHB003 following. However, the HFM is not for the moment on the priority list. 

## Marta's notes

- **UPDATE on SM18 (COVID-19 point of view):**

`Gerard` _Everything moves smoothly, especially confined to your office. The control room is the busiest place some times and must be careful when people passed by. However, distances are thoroughly respected and PPEs are used._

`Franco` _Nice to be back! It is a bit strange to work with the PPEs but everything is moving smoothly. However, the interventions during a test might be tricky, increasing the risk since no more physical space is available and the distances can be violated._ 


- **NEWS FROM SM18:**

    - Update your EDH status weekly or whenever it needed but always in time.
    -  FTEC: New arrival expected around July, filling Paula's position after she leaves.
    -  TE department meeting: discussion on all relevant aspects.
        - Special session HL-LHC: a report from SM18 will be given (Marta)
    - Engineers, Operators, Technicians on site:
        - `ENG` Franco, Gerard, Michal on-site (%).
        - `OPR`  Guillaume, Jerome, Olivier, Vincent, Maryline, Gaelle on-site (%). 
        - `TEC` Bertrand and Abdhelay on-site (%). Raphael will also start next week.

    -  Posters will be placed at SM18 in the coming days.

    -  Face-shields have been distributed on-site. Guillaume is participating actively in their assembly. 

- **NEWS in general:**
    1. Administrative actions per individual: No other administrative update(s) for this week.

    2. **EDH:** 

            I. REQUEST: arrange it accordingly every time for all the proper reasons:
            - Information, tracking, insurance, payment, etc.
            - Count/expected number of people on site: management monitoring the smooth transition.
            - It is not a lack of confidence, but about the legal coverage.

            II. APPROVAL: Follows the "1-to-1" rule! 
            ie., if you request a day, a week, etc. then you have to ask it a day, 
            a week, and so on in advance if possible.

    3. Follow the predefined from the risk analysis **measures** (mask, wash/clean, distance and more PPEs if needed). Anyhow, they can change or force applied if needed.

        **In brief**, online mandatory course “COVID-19 - Health & Safety Measures at CERN". PPE, masks, e-learning, posters, factsheets for different professional occasions, basic rules (personal responsibilities, distance, masks, cleaning and washing, extra PPEs, mitigation measures (transportation, common spaces: rooms, corridors, stairs, elevators, toilets, etc. Common tools, printers, water tanks, etc.) 

    4. Fill the excel file concerning the vacation periods that you probably request (file@matest). As a general recommendation, try to plan your holidays in advance! There is already discussed, not finalized [!!!], idea of _"**holiday donation**"_ for the upcoming working period, after September. 


**MORE:**

1. **SLM meeting**: 
No meeting scheduled for this week, hence, there are no other administrative updates.

2. Send your input to Marta/Gerard about your current tasks for the current week.

3. **Marta's agenda** 
Several meetings, such as the SLM, the SC-links, the IT-Strings, the construction documentation, and presentations, with the coordination of the magnet transportation. Moreover, the cryogenics meeting concerning the modifications for the test benches. Lastly, meeting for the new visitor's area and the preparation of the updated document.

    **NOTE**: Check weekly the updated version of the "_LOGBOOK Marta TF_". 


#### Documentation:

* The documentation is a "proof of work", especially for those that partially working. However, it accounts for everybody!

    **To whom it might concern**: Please provide the documentation requested by Marta. 


#### OBLIGATORY COURSE (new)

* Online course “COVID-19 - Health & Safety Measures at CERN” which is now available on the Safety Training catalog on CERN’s Learning Hub (https://learninghub.cern.ch) > Catalog: Safety > Domain: Safety and Health (SH).


## Around the table discussion

Here is presented a brief summary of everyone's entries:

**Bernardo**: 
- Working on the collection of the data from PSI. 
- Meeting with Erica. 
- Continuing with an online training course (duration: 7 weeks, 3x/week)


**Stathis**: 
- Continuing the development of the automated word report for the quadrupole diodes: template. 
- Continuing the development of the additional functionality for the users (File manipulation, Copy/Paste, Pre-processing, and values exportation from Diadem to Excel, create from scratch the diodes calibration files form the initial test files, and more).
- Preparation of the minutes of the TW meeting.
- Meetings attention: supervisor, scientific, Masterclass.
- Finalizing the yearly affiliation report from the University's committee.


**Franco**: 
- Meetings/Discussions about magnet test planning and confirmation.
- Feather 2 test preliminary discussions.
- Test with Vincent on the 11T.
- Upcoming planned tests:
    - PC test: 700V above ground: Simulate energy extraction in magnet during a quench in "real" machine condition (0.5-1 second maximum duration). 
    - Intermediate shunt between the apertures test. Preliminary discussions and simulation studies.


**Gaelle**: 
- Attended some meetings such as the MBHA meeting, for the problem of the segment position.
- Diadem training (ongoing)
- Patroling of SM18
- QXF magnet preparations on site


**Gerard**: 
- Attended several meetings, like the MBHA-001 meeting, MP3, HTS and more
- Meetings with Supervisees
-  Update on 11T (with Franco, Gaelle, Vincent, and ELQA team)
    - Two unexpected quenches occurred, indicating degradation in the coil... Seems to be serious and it is under investigation
- At the same time : Special "voltage to ground" test (1kV from the ground) is under discussion and preparation in order to study: 
    - a. spikes on magnet and 
    - b. voltage relation between Quench Heaters to coil connections.


**Guillaume**: 
- Works on-site at SM18.


**Michal**: 

- Cluster D Updates (with Jerome, Guillaume): 
     Resumed last week, with the magnet being at 4.5K. The electronics have been restarted except the power converter... which is still oscillating and triggers the QPS. Several parameters have been tested, such as the inductance but everything seems to be fine. Additionally, several resets have been addressed to the PLCs, on the switches but nothing seems to affect the PC's behavior.
    - Scheduled Plans: VI measurement at 4.5K for each aperture, in order to check if there is any degradation after the cooldown.
 
- Updates on DEMO2: 
    - The pressure test has been performed and it is fine (Amalia). Cables are ready, with 2 racks being ready from Jerome, Guillaume. 
  - The Water-cooled cable team is expected at SM18 today... but not yet arrived.
 - Scaffolding is ready (Bertrand, Abdhelay)
 - Cooling the feedbox tomorrow (Jerome Fl.)

- Updates on S-Links: 
    Organization of the test (maximum 3-5 days), which can be performed in parallel with the training of 11T.


**Ioannis**: 
- Carpenter V2 development: specifically on the interactive wizard feature. It will be presented on the 2nd Scientific Meeting of the section!
In general, everything is going according to the plan!
- Database changes discussions
- Attended several meetings/MasterClass
- Carpenter meeting on Reporting


**Maryline**:
- Works on-site at SM18 this week.


**Olivier**: 
- Resume work on-site at SM18 for the test of the QXFP1.


**Patrick**:
- Patroling of the SM18
- SM18 safety preparations.


**Salvador**:
- Preparation of the DEMO2 planning 
- I attended several meetings and prepared the minutes of several of them.
- Start the electrical layout of the cluster F, while it is pending input.


**Yacine**:
- 2kA cable test safety assessment document.
- Meeting with Cryogenics concerning the 2kA cable test.
- A1 and F1: The preliminary version is finished. Comments have been sent to Yacine. He will continue working on them.
- Attended several meetings


**Paula**:
- Integration meeting prep and minutes (services).
- Provided feedback on several other documents.
- Instruction documents preparation (+sharepoint management).
- Attended several meetings, such as the Power converter installation meeting, the DFX at the tunnel meeting, and more.
- Contribution documents: SC-links, IT String, Transportation, DFX at the tunnel.
- Visitor service work


**Jerome**: 
- Cluster D modification meeting
- Cluster F meeting with Maryline
- Cryogenics meeting
- Working with Michal on the magnet test cabling


**Vincent**:
- Works on the field at SM18 for the test of the QXFP1, and 11T.


**Ruben**:
- Continuing on the simulation results and expanding his model by the introduction of different types of defects.

-------

There are minutes attached to this event. Show them.
The agenda of this meeting is empty